# Dimitrios Kostas [^1] Mathematics subject classification:65M60,65M12,35K55

### v2

@@ 1 Introduction
The complex Landau equation, a fundamental model in Nonlinear dynamics, has received limited attention in the existing literature regarding its nonlinear dynamics, with only a handful of studies addressing this topic.To the best of our knowledge, the application of Discontinuous Galerkin (DG) methods—specifically the Symmetric (SIPG), Nonsymmetric (NIPG), and Incomplete (IIPG) Interior Penalty Galerkin methods—to the complex Landau equation remains unexplored.This study addresses the critical gap by introducing and analyzing these three Discontinuous Galerkin formulations for the complex Landau equation. Our work establishes novel stability estimates for the proposed schemes and demonstrates their convergence behavior through detailed numerical and theoretical analysis.
@@
In this section, we review key results that form the theoretical and computational foundation of our work. he Discontinuous Galerkin (DG) framework has become an established tool in computational mathematics, frequently employed due to its favorable convergence properties and capacity to maintain low numerical errors across diverse problem classes. The fundamental theoretical and computational basis for DG methods is well-documented. A rigorous and comprehensive theoretical foundation, including the Symmetric Interior Penalty Galerkin (SIPG) formulation, is provided by Di Pietro and Ern [7], which also offers detailed stability analyses for both elliptic and hyperbolic partial differential equations. Further foundational insight is provided by the monograph of Cockburn et al. [6], which furnishes a detailed account of the mathematical principles underpinning interior penalty DG methods, specifically addressing stability, consistency, error estimation, and approximation properties.
@@
Advancements in the DG methodology include extensions to complex physical systems. A shock-capturing variant of SIPG, specifically tailored to convection–diffusion–reaction systems characterized by steep gradients or discontinuities, is discussed by Yücel and Karasözen [25]. For practical, time-dependent implementations, Akman et al. [4] demonstrate the effectiveness of SIPG when applied to optimal control problems governed by unsteady convection–diffusion equations. The efficiency and scalability of high-order DG discretizations have been addressed through the development of hybrid multigrid preconditioners for the incompressible Navier–Stokes equations by Feng and Wu [9]. Complementarily, the combination of DG discretization with a reduced basis approach has been explored for fluid dynamics problems with geometrically parametrized domains, yielding substantial computational savings via model order reduction [23].
@@
The foundational formulation of the interior penalty (IP) method utilizing discontinuous elements was established by Arnold in his seminal work [3], providing the theoretical genesis for subsequent developments, including the widely adopted SIPG, Nonsymmetric (NIPG), and Incomplete (IIPG) Interior Penalty Galerkin schemes. The flexibility of DG methods in handling complex domains is further evidenced by its successful application in non-overlapping domain decomposition strategies for the incompressible Navier–Stokes equations [12]. Furthermore, efforts to ensure computational reliability and efficiency have driven the development of A-posteriori error estimation techniques for DG discretizations of elliptic problems [16], culminating in robust adaptive frameworks that offer convergence guarantees [17]. In the context of computational efficiency, a simplified DG scheme for diffusion-dominated problems has also been proposed, which emphasizes robustness without compromising accuracy [24]. The robustness and accuracy of discontinuous hp-Finite Element Methods in resolving sharp layers in reaction diffusion problems, particularly in high-Péclet number regimes, is thoroughly investigated in Houston et al. [14].
@@
While the present work is centered on the complex Landau equation, related literature provides structural and numerical context. For instance, the multiscale framework for solving the Ginzburg–Landau equation [5] offers relevant insights for DG-based multiscale approaches to nonlinear PDEs, despite its focus on vortex-capturing spaces. Similarly, the Crank–Nicolson scheme applied to the Schrödinger equation [11] shares a structural similarity with the complex Landau equation—specifically the presence of imaginary terms—yet the underlying mathematical structures and physical applications differ significantly. We note that DG-based schemes have been studied for the nonlinear Fokker–Planck–Landau equation [10, 13, 15]; however, these schemes diverge considerably from the formulation introduced here, both in terms of the underlying numerical methods and the intended physical applications. Furthermore, studies examining nonlinear systems that address temporal, rather than spatial, discontinuities [19, 20] do not involve the essential feature of complex-valued solutions and imaginary components inherent to the complex Landau equation. Finally, Petrov–Galerkin finite element methods [22, 26] represent an alternative approach to DG methods. Nonetheless, certain stabilization techniques developed in the Petrov–Galerkin context may offer transferable insights for handling discontinuities and ensuring stability within the DG framework considered in this study. Related conceptual ideas concerning stabilization and discretization are also found in the study of elliptic partial differential equations by Rivière et al. [18].
@@ 2.1 The Landau equation
$\frac{du}{dx}=u+(1+ia)\Delta u-(1+ib)|u|^{2}u$ (2.2)
In above equation $a,b\in\Re$ are real parameters that significantly influence the system’s temporal and spatial behavior. The presece of the imaginary component in the diffusion term $(1+ib)|u|^{2}u$ gives rise to nonlinear frequency modulation, in which the wave frequency becomes ampiltude-dependent.
The complex coefficients play a crucial role in determining the qualitive nature of the solutions. In particular, they lead to a variety of rich spatiotemporial including
@@
Let $T_{h}$ the mesh of (Triangles/quads) that we defined and $E_{h}$ the set of interrior edges in 2D or faces in 3D and $u_{h}$ a function of the space $V_{h}$, $u_{h}\in V_{h}$
@@
Grönwall inequality-Grönwall lemma
Suppose a nonegative sequence $\{a_{n}\}$ satisfy
$a_{n}\leq aa_{n-1}+C\Delta t\ \ \ n=1,2$... (2.10)
where
  • $a_{0}\geq 0$ is known,
  • $a\geq 0$ is a constant (often of the form $a=1+C\Delta t$
  • The constant $C\geq 0$
  • $\Delta t>0$ the time step
@@
$a_{N}\leq a_{0}a^{N}+C\Delta t\sum_{j=0}^{N-1}a^{j}=a_{0}a^{N}+C\Delta t\frac{a^{N}-1}{a-1},ifa\neq 1$ (2.11)
If now a=1 the the inequality simpified to
$a_{N}\leq a_{0}+C\Delta t$
@@
where $u_{h},v_{h}\in\text{V}_{h}$ The belinear form $a_{h}(u_{h},v_{h})$ was firstly introduced by [3] and represents the standard weak formulation of the Laplacian operator, where $v_{h}$ denotes the test function.
The formulation depends on the parameter $\theta$, leading to the following variants of Interrior Penalty Galerkin method
@@
The part b ios often assumed to be non-negative to avoid unphysical behavior b¿0.
@@
Find $(u_{h1},u_{h2})\in V_{h}\times V_{h}$ sich that for all $(v_{h1},v_{h2})\in V_{h}\times V_{h}$
@@ 2.3 Existence and Uniqueness of the Continuous and Discrete Problems
The Landau equation, in its regularized form, is a nonlinear parabolic partial differential equation (PDE) of Fokker–Planck type. Under standard assumptions on the regularity and decay of the initial data, and with appropriate smoothing of the collision kernel, classical results guarantee the existence and uniqueness of weak solutions in suitable spaces $L^{2}(0,T,H^{1}(\Omega))$ Reference [1] Discusses the ,derivation, justification, and weak solution of the Landau equation from physical models. A comprehensive of well-possessiveness and convergence results for discretized (numerical) versions of the Landau equation is provided in [11], while [21] establishes results for the linearized Landau operator.
The fully discrete problem, where we approximate the Landau equation using a discontinuous Galerkin (DG) or continuous Galerkin (CG) finite element scheme, we consider the discrete formulation in a finite-dimensional function space. Due to the finite dimensionality and the local Lipschitz continuity of the nonlinear discrete operator, the discrete problem admits a unique solution that is local in time . Furthermore, under standard coercivity and boundedness assumptions on the discrete bi-linear forms—particularly for DG schemes equipped with appropriate penalty parameters—energy estimates can be derived to ensure global-in-time existence and uniqueness of the discrete solution. A brief stability analysis further confirms that the numerical method is well-posed and robust under mesh refinement.
@@ 3 Stability Estimate
Step 3 Estimate the right-hand side using inequalities such as Young’s inequality and inverse inequalities appropriate for DG spaces.
Step 4 Absorb the resulting terms into the left-hand side of the inequality using coercivity of the bilinear form.
Step 5 Apply Grönwall’s lemma to conclude a discrete energy estimate and establish stability.
@@
Let $\Omega\subset\Re^{d}$be a bounded Lipschitz domain, and ${T}^{n}_{i=1}$ be a comforming shape regular partition of $\Omega$ into the disjoint element such that $\Omega=\cup_{i=1}^{N}T^{i}$ with $T_{i}\cap T_{j}=$ for $i\neq j$ and time interval be discretized into a partition in time $0=t^{1}<t^{2}<...<t^{N}=T.$ Assume that the bilinear form $a_{h}(u_{h},v),$ defined in equation 2.13, corresponds to a time discretization using theta scheme ,and a special discretization based on one of the interrior penalty discontinuous Galerkin methods (SIPG),(NIPG),(IIPG) For penalty parameters $\sigma$ large enough If it holds the 2.12 Poincare’ inequality. Then there exists constants $\beta>0,c\geq 0$ and $C>0$ independent of the mesh size h and time step $\Delta t$, that the bi-linear term satisfies the following estimates:
@@
Now adding the above equations 3.20 and 3.21 and rearrange the terms by moving them to the left-hand side or the opposite and using coercivity to absorb them so we obtain
@@
Now using the inequality
[…]
$\Delta tmax(\|u_{h1}\|^{2}_{L^{2}(\Omega)}+\|u_{h2}\|^{2}_{L^{2}(\Omega)})\leq\int_{\Omega}(u_{h1}^{2}+u^{2}_{h2})^{2}dxdt$
The above hold since
$\Delta tmax(\|u_{h1}\|^{2}_{L^{2}(\Omega)}+\|u_{h2}\|^{2}_{L^{2}(\Omega)})\leq 0$
and
@@
$C(\frac{\|u^{n-1}_{h1}\|^{2}_{L^{2}(\Omega)}}{2}+\frac{\|u^{n-1}_{h2}\|^{2}_{L^{2}(\Omega)}}{2})+c\Delta t$ (3.23)
Now we obntain the scheme
$a^{n}\leq Ca^{n-1}+c\Delta t$
. where $a^{n}=\frac{\|u^{n}_{h1}\|^{2}_{L^{2}(\Omega)}}{2}+\frac{\|u^{n}_{h2}\|^{2}_{L^{2}(\Omega)}}{2},c(max(\|u_{h1}\|^{2}_{DG}+\|u_{h2}\|^{2}_{DG}),max(\|u_{h1}\|^{2}+\|u_{h2}\|^{2}))$.
Using Grönwall inequality we obtain the result $a^{N}\leq a^{0}C^{N}+c\Delta t\frac{C^{N}-1}{C-1}$ which is the desirable result when $C\neq 1$
If $C=1$ then $a^{N}\leq a^{0}+cN\Delta t$

### v3

@@ 1 Introduction
The complex Landau equation, a fundamental model in Nonlinear dynamics, has received limited attention in the existing literature regarding its nonlinear dynamics, with only a handful of studies addressing this topic.To the best of our knowledge, the application of Discontinuous Galerkin (DG) methods-specifically the Symmetric (SIPG), Nonsymmetric (NIPG), and Incomplete (IIPG) Interior Penalty Galerkin methods to the complex Landau equation remains unexplored.This study addresses the critical gap by introducing and analyzing these three Discontinuous Galerkin formulations for the complex Landau equation. Our work establishes novel stability estimates for the proposed schemes and demonstrates their convergence behavior through detailed numerical and theoretical analysis.
@@
In this section, we review key results that form the theoretical and computational foundation of our work. The Discontinuous Galerkin (DG) framework has become an established tool in computational mathematics, frequently employed due to its favorable convergence properties and capacity to maintain low numerical errors across diverse problem classes. The fundamental theoretical and computational basis for DG methods is well-documented. A rigorous and comprehensive theoretical foundation, including the Symmetric Interior Penalty Galerkin (SIPG) formulation, is provided by Di Pietro and Ern [9], which also offers detailed stability analyses for both elliptic and hyperbolic partial differential equations. Further foundational insight is provided by the monograph of Cockburn et al. [8], which furnishes a detailed account of the mathematical principles underpinning interior penalty DG methods, specifically addressing stability, consistency, error estimation, and approximation properties.
@@
Advancements in the DG methodology include extensions to complex physical systems. For practical, time-dependent implementations, Akman et al. [4] demonstrate the effectiveness of SIPG when applied to optimal control problems governed by unsteady convection diffusion equations. The efficiency and scalability of high-order DG discretizations have been addressed through the development of hybrid multigrid preconditioners for the incompressible Navier-Stokes equations. Complementarily, the combination of DG discretization with a reduced basis approach has been explored for fluid dynamics problems with geometrically parametrized domains, yielding substantial computational savings via model order reduction [22].
@@
The foundational formulation of the interior penalty (IP) method utilizing discontinuous elements was established by Arnold in his seminal work [3], providing the theoretical genesis for subsequent developments, including the widely adopted SIPG, Nonsymmetric (NIPG), and Incomplete (IIPG) Interior Penalty Galerkin schemes. The flexibility of DG methods in handling complex domains is further evidenced by its successful application in non-overlapping domain decomposition strategies for the incompressible Navier-Stokes equations [13]. Furthermore, efforts to ensure computational reliability and efficiency have driven the development of A-posteriori error estimation techniques for DG discretizations of elliptic problems [17], culminating in robust adaptive frameworks that offer convergence guarantees [18]. In the context of computational efficiency, a simplified DG scheme for diffusion-dominated problems has also been proposed, which emphasizes robustness without compromising accuracy [23]. The robustness and accuracy of discontinuous hp-Finite Element Methods in resolving sharp layers in reaction diffusion problems, particularly in high-Peclet number regimes, is thoroughly investigated in Houston et al. [15].
@@
While the present work is centered on the complex Landau equation, related literature provides structural and numerical context. For instance, the multiscale framework for solving the Ginzburg-Landau equation [5] offers relevant insights for DG-based multiscale approaches to nonlinear PDEs, despite its focus on vortex-capturing spaces. Similarly, the Crank-Nicolson scheme applied to the Schrödinger equation [12] shares a structural similarity with the complex Landau equation specifically the presence of imaginary terms yet the underlying mathematical structures and physical applications differ significantly. We note that DG-based schemes have been studied for the nonlinear Fokker-Planck-Landau equation [11, 14, 16]; however, these schemes diverge considerably from the formulation introduced here, both in terms of the underlying numerical methods and the intended physical applications. Furthermore, studies examining nonlinear systems that address temporal, rather than spatial, discontinuities [6, 7] do not involve the essential feature of complex-valued solutions and imaginary components inherent to the complex Landau equation. Finally, Petrov-Galerkin finite element methods [21, 26] represent an alternative approach to DG methods. Nonetheless, certain stabilization techniques developed in the Petrov-Galerkin context may offer transferable insights for handling discontinuities and ensuring stability within the DG framework considered in this study. Related conceptual ideas concerning stabilization and discretization are also found in the study of elliptic partial differential equations by Riviëre et al. [19].
@@ 2.1 The Landau equation
$\frac{\partial u}{\partial t}=u+(1+ia)\Delta u-(1+ib)|u|^{2}u$ (2.2)
In above equation $a,b\in\Re$ are real parameters that significantly influence the system’s temporal and spatial behavior. The presence of the imaginary component in the diffusion term $(1+ib)|u|^{2}u$ gives rise to nonlinear frequency modulation, in which the wave frequency becomes amplitude-dependent.
The complex coefficients play a crucial role in determining the qualitive nature of the solutions. In particular, they lead to a variety of rich spatiotemporal phenomena including
@@
Let $T_{h}$ the mesh of (Triangles/quads) that we defined and $E_{h}$ the set of interior edges in 2D or faces in 3D and $u_{h}$ a function of the space $V_{h}$, $u_{h}\in V_{h}$
@@
Gronwall inequality-Gronwall lemma
[…]
$a_{n}\leq e^{c\Delta t}a_{n-1}\ \ \ n=1,2$... (2.10)
@@
$a_{N}\leq e^{cN\Delta t}a_{0}$ (2.11)
@@
where $u_{h},v_{h}\in\text{V}_{h}$ The bilinear form $a_{h}(u_{h},v_{h})$ was firstly introduced by [3] and represents the standard weak formulation of the Laplacian operator, where $v_{h}$ denotes the test function.
The formulation depends on the parameter $\theta$, leading to the following variants of Interior Penalty Galerkin method
@@
The part b is often assumed to be non-negative to avoid unphysical behavior b¿0.
@@
Find $(u_{h1},u_{h2})\in V_{h}\times V_{h}$ such that for all $(v_{h1},v_{h2})\in V_{h}\times V_{h}$
@@ 2.3 Existence and Uniqueness of the Continuous and Discrete Problems
The Landau equation, in its regularized form, is a nonlinear parabolic partial differential equation (PDE) of Fokker-Planck type. Under standard assumptions on the regularity and decay of the initial data, and with appropriate smoothing of the collision kernel, classical results guarantee the existence and uniqueness of weak solutions in suitable spaces $L^{2}(0,T,H^{1}(\Omega))$ Reference [1] Discusses the ,derivation, justification, and weak solution of the Landau equation from physical models. A comprehensive of well-possessiveness and convergence results for discretized (numerical) versions of the Landau equation is provided in [12], while [20] establishes results for the linearized Landau operator.
The fully discrete problem, where we approximate the Landau equation using a discontinuous Galerkin (DG) or continuous Galerkin (CG) finite element scheme, we consider the discrete formulation in a finite-dimensional function space. Due to the finite dimensionality and the local Lipschitz continuity of the nonlinear discrete operator, the discrete problem admits a unique solution that is local in time . Furthermore, under standard coercivity and boundedness assumptions on the discrete bi-linear forms-particularly for DG schemes equipped with appropriate penalty parameters-energy estimates can be derived to ensure global-in-time existence and uniqueness of the discrete solution. A brief stability analysis further confirms that the numerical method is well-posed and robust under mesh refinement.
@@ 3 Stability Estimate
Step 3 Absorb the resulting terms into the left-hand side of the inequality using coercivity of the bilinear form.
Step 4 Apply Gronwall lemma to conclude a discrete energy estimate and establish stability.
@@
Let $\Omega\subset\Re^{d}$be a bounded Lipschitz domain, and ${T}^{n}_{i=1}$ be a conforming shape regular partition of $\Omega$ into the disjoint element such that $\Omega=\cup_{i=1}^{N}T^{i}$ with $T_{i}\cap T_{j}=$ for $i\neq j$ and time interval be discretized into a partition in time $0=t^{1}<t^{2}<...<t^{N}=T.$ Assume that the bilinear form $a_{h}(u_{h},v),$ defined in equation 2.13, corresponds to a time discretization using theta scheme ,and a special discretization based on one of the interior penalty discontinuous Galerkin methods (SIPG),(NIPG),(IIPG) For penalty parameters $\sigma$ large enough If it holds the 2.12 Poincare inequality. Then there exists constants $\beta>0,c\geq 0$ and $C>0$ independent of the mesh size h and time step $\Delta t$, that the bi-linear term satisfies the following estimates:
@@
Lemma 3.3 (A Priori $L^{\infty}(L^{2})$ Stability of semi discrete system ).
Let $(u_{h1},u_{h2})$ be the solution to the semi-discrete system, and the $a(u,v)$ symmetric since our method is Symmetric Interior Penalty Galerkin (SIPG).Then for all $t\in[t^{n-1},t^{n}]$ the following bound holds:
$\|u_{h1}(t)\|^{2}_{L^{2}(\Omega)}+\|u_{h2}(t)\|^{2}_{L^{2}(\Omega)}\leq\left(\|u_{h1}(t^{n-1})\|^{2}_{L^{2}(\Omega)}+\|u_{h2}(t^{n-1})\|^{2}_{L^{2}(\Omega)}\right)e^{2\Delta t}.$ (3.20)
Consequently, $(u_{h1},u_{h2})\in L^{\infty}(t^{n-1},t^{n};L^{2}(\Omega))$.
Proof.
Testing the semi-discrete system with $(u_{h1},u_{h2})$ and summing the resulting equations, we obtain the differential form of Equation 3. By dropping the quadratic non-negative terms $a_{h}(u_{hi},u_{hi}),i=1,2$ and the dissipative quartic term $\int_{\Omega}(u_{h1}^{2}+u_{h2}^{2})^{2}dx$, we arrive at:
[…]
Applying the integral form of Gronwall’s inequality over the interval $t^{n-1},t]$ for any $t\leq t^{n}$ yields the desired exponential bound. Since the right-hand side is finite and independent of $t$, the solution is bounded in $L^{\infty}(t^{n-1},t^{n},L^{2}(\Omega))$. ∎
Remark 3.4.
From previous 3.3 , integrating in $t^{n-1},t^{n}]$ we obtain
$\int_{t^{n-1}}^{t^{n}}\Big(\|u_{h1}(t)\|^{2}_{L^{2}(\Omega)}+\|u_{h2}(t)\|^{2}_{L^{2}(\Omega)}\Big)dt\leq\int_{t^{n-1}}^{t^{n}}\Big(\|u_{h1}(t^{n-1})\|^{2}_{L^{2}(\Omega)}+\|u_{h2}(t^{n-1})\|^{2}_{L^{2}(\Omega)}\Big)e^{2(t-t^{n-1})}dt$ (3.21)
or equivalently
$\int_{t^{n-1}}^{t^{n}}\Big(\|u_{h1}(t)\|^{2}_{L^{2}(\Omega)}+\|u_{h2}(t)\|^{2}_{L^{2}(\Omega)}\Big)dt\leq\Big(\|u_{h1}(t^{n-1})\|^{2}_{L^{2}(\Omega)}+\|u_{h2}(t^{n-1})\|^{2}_{L^{2}(\Omega)}\Big)\frac{e^{2\Delta t}-1}{2}$ (3.22)
Now from Taylor $\frac{e^{2\Delta t}-1}{2}\approx\Delta t+O(\Delta t^{2})$ we can confirm that the integral is bounded.
@@
$\|u_{h1}^{n}\|^{2}_{L^{2}}+\|u_{h2}^{n}\|^{2}_{L^{2}}\leq e^{CT}\left(\|u_{h1}^{0}\|^{2}_{L^{2}}+\|u_{h2}^{0}\|^{2}_{L^{2}}\right),$ (3.23)
where $T=N\Delta t$. This implies the scheme is stable in $L^{\infty}(0,T;L^{2}(\Omega))$.
@@
Now adding the above equations 3.24 and 3.25 and rearrange the terms by moving them to the left-hand side or the opposite and using coercivity to absorb them
  • In case of SIPG we obtain
@@
    By Remark 3.3, the solution is bounded in $L^{\infty}(t^{n-1},t^{n};L^{2}(\Omega))$. Therefore, from remark 3.4 the integral on the right-hand side is well-defined and can be bounded : where $\Delta t=t^{n}-t^{n-1}$ Also we observe that
@@
    $e^{2\Delta t}(\frac{\|u^{n-1}_{h1}\|^{2}_{L^{2}(\Omega)}}{2}+\frac{\|u^{n-1}_{h2}\|^{2}_{L^{2}(\Omega)}}{2})$ (3.27)
    Using the coercivity of the bilinear form $a_{h}(\cdot,\cdot)$, there exists a constant $\alpha>0$ such that:
    $\int_{t^{n-1}}^{t^{n}}a_{h}(u_{h1},u_{h1})+a_{h}(u_{h2},u_{h2})dt\geq\alpha\Delta t\Big(\|u_{h1}\|^{2}_{DG}+\|u_{h2}\|^{2}_{DG}\Big)\geq 0$
    $\frac{\|u^{n}_{h1}\|^{2}_{L^{2}(\Omega)}}{2}+\frac{\|u^{n}_{h2}\|^{2}_{L^{2}(\Omega)}}{2}\leq e^{2\Delta t}(\frac{\|u^{n-1}_{h1}\|^{2}_{L^{2}(\Omega)}}{2}+\frac{\|u^{n-1}_{h2}\|^{2}_{L^{2}(\Omega)}}{2})$
    Now we obtain the scheme
    $a^{n}\leq e^{2\Delta t}a^{n-1}$
    . where $a^{n}=\frac{\|u^{n}_{h1}\|^{2}_{L^{2}(\Omega)}}{2}+\frac{\|u^{n}_{h2}\|^{2}_{L^{2}(\Omega)}}{2}$.
    Using Grönwall inequality we obtain the result
    $a^{N}\leq e^{2N\Delta t}a^{0}$ (3.28)
    which is the desirable result.
  • In case of IIPG or NIPG we use that
    $\int_{t^{n-1}}^{t^{n}}a_{h}(u_{h1},u_{h1})+a_{h}(u_{h2},u_{h2})dt\geq\int_{t^{n-1}}^{t^{n}}c(\|u_{h1}\|^{2}_{DG}+\|u_{h2}\|^{2}_{DG})dt,c>0$
    In addition we have
    $|a_{h}(u_{h1},u_{h2})|\leq Mc\|u_{h1}\|_{DG}\|u_{h2}\|_{DG}\leq\frac{\gamma}{2}(\|u_{h1}\|^{2}_{DG}+\|u_{h2}\|^{2}_{DG})$
    where M comes from continuity of $a_{h}(.,.)$ and $\gamma\geq Mc$, $\gamma$ is a constant that depends on:The shape regularity of the mesh.The polynomial degree $p$ of your basis functions (usually $\gamma\propto p^{2}$).The coupling coefficients . From3.25 we obtain
    $\frac{\|u^{n}_{h1}\|^{2}_{L^{2}(\Omega)}}{2}+\frac{\|u^{n}_{h2}\|^{2}_{L^{2}(\Omega)}}{2}+\int_{t^{n-1}}^{t^{n}}c(\|u_{h1}\|^{2}_{DG}+\|u_{h2}\|^{2}_{DG})+\int_{\Omega}(u_{h1}^{2}+u^{2}_{h2})^{2}dxdt\leq$
    $\frac{\|u^{n-1}_{h1}\|^{2}_{L^{2}(\Omega)}}{2}+\frac{\|u^{n-1}_{h2}\|^{2}_{L^{2}(\Omega)}}{2}+\int_{t^{n-1}}^{t^{n}}\|u_{h1}\|^{2}_{L^{2}(\Omega)}+\|u_{h2}\|^{2}_{L^{2}(\Omega)}+\frac{\gamma}{2}(\|u_{h1}\|^{2}_{DG}+\|u_{h2}\|^{2}_{DG})dt$
    or equivalently
    $\frac{\|u^{n}_{h1}\|^{2}_{L^{2}(\Omega)}}{2}+\frac{\|u^{n}_{h2}\|^{2}_{L^{2}(\Omega)}}{2}+(c-\frac{\gamma}{2})\int_{t^{n-1}}^{t^{n}}(\|u_{h1}\|^{2}_{DG}+\|u_{h2}\|^{2}_{DG})dt+\int_{\Omega}(u_{h1}^{2}+u^{2}_{h2})^{2}dxdt\leq$
    $\frac{\|u^{n-1}_{h1}\|^{2}_{L^{2}(\Omega)}}{2}+\frac{\|u^{n-1}_{h2}\|^{2}_{L^{2}(\Omega)}}{2}+\int_{t^{n-1}}^{t^{n}}\|u_{h1}\|^{2}_{L^{2}(\Omega)}+\|u_{h2}\|^{2}_{L^{2}(\Omega)}dt$
    where it holds that $c-\frac{\gamma}{2}>0$ since these constants depend on the shape regularity of the mesh and the penalty parameter $\sigma$ and $(c-\frac{\gamma}{2})\int_{t^{n-1}}^{t^{n}}(\|u_{h1}\|^{2}_{DG}+\|u_{h2}\|^{2}_{DG})dt+\int_{\Omega}(u_{h1}^{2}+u^{2}_{h2})^{2}dxdt>0$ from here we obtain the scheme
    $\frac{a^{n}}{2}\leq\frac{a^{n-1}}{2}+\int_{t^{n-1}}^{t^{n}}a(t)dt$
    where
    $\frac{a^{n}}{2}=\frac{\|u^{n}_{h1}\|^{2}_{L^{2}(\Omega)}}{2}+\frac{\|u^{n}_{h2}\|^{2}_{L^{2}(\Omega)}}{2}$
    $\frac{a^{n-1}}{2}=\frac{\|u^{n-1}_{h1}\|^{2}_{L^{2}(\Omega)}}{2}+\frac{\|u^{n-1}_{h2}\|^{2}_{L^{2}(\Omega)}}{2}$
    $\int_{t^{n-1}}^{t^{n}}a(t)dt=\int_{t^{n-1}}^{t^{n}}\|u_{h1}\|^{2}_{L^{2}(\Omega)}+\|u_{h2}\|^{2}_{L^{2}(\Omega)}dt$
    Finally after using Grönwall inequality the desirable result 3. As long as $\sigma$ is chosen large enough (for IIPG) and $\gamma$ is a fixed physical constant of the system, the stability is uniform.
Remark 3.6.
The theoretical lower bound for the penalty parameter is typically expressed as:
$\sigma>\sigma_{0}=C_{inv}\cdot\frac{(p+1)(p+d)}{d}$
[…]
$\sigma>\frac{(p+1)(p+2)}{2}$
If $p=1$ (linear elements), $\sigma$ should be at least $3$. For $p=2$ (quadratics), $\sigma$ should be at least $6$. In practice, most researchers double these values (e.g., using $\sigma=10$ or $20$) to ensure safety across distorted meshes.
Remark 3.7.
The stability result holds for interior penalty discontinuous Galerkin methods of arbitrary polynomial degree $k\geq 0$. In particular, for SIPG, NIPG, and IIPG discretizations, the DG bilinear form $a_{h}(\cdot,\cdot)$ is coercive provided that the penalty parameter $\sigma$ is chosen sufficiently large, possibly depending on the polynomial degree $k$ and the mesh shape-regularity. Under this assumption, the stability estimate remains valid for high-order DG schemes, with constants independent of the mesh size $h$ and time step $\Delta t$, but possibly depending on $k$.